OpenStudy (anonymous):

2/3x = 6 (1 point) x = 9 x = 4 x = 18 x = 12

OpenStudy (anonymous):

Alright, But How Would You Do This One? x/4 = 12/6 (1 point) x = 6 x = 12 x = 8 x = 48

OpenStudy (anonymous):

cross multiply